A person driving a car suddenly applies the brakes. The car takes 4 seconds to come to rest while traveling 20 meters at constant acceleration. Can the speed of the car immediately before the brakes were applied be determined without first determining the car's acceleration?
1
Expert's answer
2017-09-11T07:39:07-0400
Yes. The speed of the car immediately before the brakes were applied is
v=at.
d=(at^2)/2→a=2d/t^2 .
v=2d/t^2 t=2d/t=2(20)/4=10 m/s.
Answer: 10 m/s.